9.) You have an investment account that started with ​$4 comma 000 10 years ago and which now has grown to ​$8 comma 000.
a. What annual rate of return have you earned​ (you have made no additional contributions to the​ account)?
b. If the investment account earns 15 % per year from now​ on, what will the​ account's value be 10 years from​ now?
a. What annual rate of return have you earned​ (you have made no additional contributions to the​ account)?
Your annual rate of return is ___%. ​(Round to two decimal​ places.)

10.) You are thinking of purchasing a house. The house costs $ 400 comma 000. You have $ 57 comma 000 in cash that you can use as a down payment on the​ house, but you need to borrow the rest of the purchase price. The bank is offering a 30​-year mortgage that requires annual payments and has an interest rate of 10 % per year. What will be your annual payment if you sign this​ mortgage?
The annual payment is ​$____ ​(Round to the nearest​ dollar.)
Answer #1

9

a

Future value = present value*(1+ rate)^time
8000 = 4000*(1+Interest rate/100)^10
Interest rate % = 7.18

b

Future value = present value*(1+ rate)^time
Future value = 8000*(1+0.15)^10
Future value = 32364.46

10

Loan amount = price-cash payment = 400000-57000=343000

PVOrdinary Annuity = C*[(1-(1+i/100)^(-n))/(i/100)]
C = Cash flow per period
i = interest rate
n = number of payments
343000= Cash Flow*((1-(1+ 10/100)^-30)/(10/100))
Cash Flow = 36385.18